### v2

# Do Larger Language Models Imply Better Generalization? A Pretraining Scaling Law for Implicit Reasoning

Abstract
Large Language Models (LLMs) have demonstrated remarkable capabilities across a wide range of tasks requiring complex reasoning. However, the effects of scaling on their reasoning abilities remain insufficiently understood. In this paper, we introduce a synthetic multihop reasoning environment designed to closely replicate the structure and distribution of real-world large-scale knowledge graphs. Our reasoning task involves completing missing edges in the graph, which requires advanced multi-hop reasoning and mimics real-world reasoning scenarios. To evaluate this, we pretrain language models (LMs) from scratch solely on triples from the incomplete graph and assess their ability to infer the missing edges. Interestingly, we observe that overparameterization can impair reasoning performance due to excessive memorization. We investigate different factors that affect this U-shaped loss curve, including graph structure, model size, and training steps. To predict the optimal model size for a specific knowledge graph, we find an empirical scaling that linearly maps the knowledge graph search entropy to the optimal model size. This work provides new insights into the relationship between scaling and reasoning in LLMs, shedding light on possible ways to optimize their performance for reasoning tasks.
1 Introduction
Large Language Models (LLMs) have demonstrated remarkable capabilities across a wide range of tasks. Recently, the reasoning capacity of LLMs has drawn a lot of attention as it is highly correlated with LLMs’ performance on many complex real-world tasks (Wei et al., 2022a; Guo et al., 2025). While the reasoning capability is usually enhanced during the post-training stage, it is reasonable to assume that LLMs have already acquired the capability during the pretraining stage, as the post-training is on a significantly smaller scale than pretraining. Some recent work has investigated the possible mechanism of LLMs learning to reason through next-token prediction pretraining (Zhu et al., 2024; Wang et al., 2024a; b). However, the impact of pretraining scaling on LLMs’ reasoning ability remains insufficiently understood.
The general scaling behavior of language models has been investigated, including the well-known exponential scaling laws for testing loss and compute proposed by Kaplan et al. (2020) and the training compute-optimal scaling studied by Hoffmann et al. (2022a). Recent work has also examined the scaling of specific capabilities like machine translation (Ghorbani et al., 2022) and knowledge capacity/memorization (Allen-Zhu & Li, 2025; Lu et al., 2024). According to these existing scaling laws, it is in general believed that larger models imply better testing loss or task performance when trained on more data.
In this paper, however, we find that the reasoning ability of language models under pretraining scaling can behave differently from normal power-law scaling, in a simplified pretraining environment. More specifically, when given enough compute, we find that the testing loss scaling curve is U-shape and there exists an optimal model size that produces the best reasoning performance/testing loss. This implies that overparameterization might hurt reasoning capability during pre-training. We first observe this phenomenon with real-world knowledge graph data, and then systematically study it through synthetically generated data.
We choose to mimic the real-world knowledge structure and distribution with synthetic knowledge graphs (KGs). We define reasoning over world knowledge as completing missing edges in an incomplete knowledge graph, which requires multiple jumps on the knowledge graph according to some pre-defined rules that are latently encoded into the graph generation process. To analyze this, we pretrain LMs from scratch using only triples from the incomplete graph and evaluate their ability to infer missing connections.
We investigate important factors that affect the U-shape scaling of reasoning loss versus language model size. Our important findings can be summarized as follows:
  • The minimum reasoning loss/maximum reasoning accuracy that a language model can reach is capped by the training data, regardless of the training steps and model size.
  • The optimal model size for a training corpus is largely fixed regardless of the training steps when the number of training steps is large enough.
  • When the underlying knowledge graph is fixed, training on more data sampled from the graph increases the optimal model size and reasoning performance.
  • More complex knowledge graph implies a larger optimal model size.
As we observed that the optimal model size is likely solely determined by the training knowledge graph, we then aim to find an empirical scaling law that can predict the optimal model size with knowledge graph statistics. We then discover a linear relationship between the optimal model size and a newly proposed graph search entropy, which measures the entropy of performing random searches on a knowledge graph. Roughly, 124 additional parameters in the optimal model size are required per 1-bit entropy increase in searching a knowledge graph.
Our work contributes to the broader understanding of LLM reasoning by shedding light on the intricate relationship between scaling and reasoning capability. Our proposed empirical reasoning scaling law provides possible practical insights for optimizing LLMs’ reasoning ability at pretraining time.
2 Preliminaries
While the real-world LLMs are pretrained on large scale text corpus, this corpus can be viewed as encoding a wide range of world knowledge. The power of LLMs lies in the fact that they can not only memorize the world knowledge and extract the knowledge when queried, but also reason over the world knowledge and draw novel conclusions. In this paper, we propose to construct a simplified pretraining corpus directly from a knowledge graph. A knowledge graph is comprised of a set of triples, and we use each knowledge triple as a training example. We test the reasoning capability of a language model trained on such a corpus by testing its accuracy in completing triples that have never been seen in the knowledge graph but can be deduced through latent rules encoded in the graph structure. For example, if we know A is B’s father, and B is C’s father, then we can deduce that A is C’s grandfather.
Formally, a knowledge graph $G$ consists of $|G|=N$ triples $(e^{h},r,e^{t})$, where $e^{h}\in\mathcal{E}$ is the head entity, $e^{t}\in\mathcal{E}$ is the tail entity, and $r\in\mathcal{R}$ is a relation. A simple example of knowledge triple is (DC, is the capital of, USA). These knowledge triples naturally form a graph, with nodes as the entities and each edge labeled with a relation type. We denote the total number of entities or nodes by $|\mathcal{E}|=N_{e}$ and the total number of edge or relation types by $|\mathcal{R}|=N_{r}$. Then a corpus constructed from this knowledge graph would consist of $N$ data points. The objective of a language model with parameter $\theta$ trained on this corpus is then:
$L(\theta)=\arg\min_{\theta}\frac{1}{N}\sum_{i=1}^{N}-\log P_{\theta}(e^{h}_{i},r_{i},e^{t}_{i}).$
To eliminate confounding variables and information contained in the lexical form of the entity and relation names, we label each entity and relation with a random ID and tokenize the IDs by characters. We use the LlaMA (Touvron et al., 2023) model architecture to implement LMs of different sizes by adjusting the hidden dimensions and the number of layers. The specific parameter scheme can be found in the Appendix A.
To evaluate the language model’s capability of reasoning over the knowledge graph, we test the language models on a held-out set of triples that are not seen in the training time. Note that all entity and relation types should have been seen during training time and the language model is only tasked to connect missing edges. To eliminate the need to generate the correct form of relation and entity IDs, and to handle the case where multiple correct answers exist, we design the testing set to be 10-option multiple-choice questions: the language model is tasked to choose the correct tail entity given the head entity and the relation. We ensure that there is only one correct answer among the given 10 options. Suppose there are $M$ questions in the testing set. [^1] We fix $M=1000$ for all of our experiments. For a ground truth triple $(e^{h},r,e^{t})$, we design 9 distracting options $e^{(1)},e^{(2)},...,e^{(9)}$. Then we use the test accuracy $\text{Acc}(\theta,G)$ and testing loss $\ell(\theta,G)$ to evaluate the reasoning capability of a language model $\theta$ over the knowledge graph $G$:
$\hat{e}_{i}$ $=\arg\max_{e\in\{e^{t}_{i},e^{(1)}_{i},e^{(2)}_{i},...,e^{(9)}_{i}\}}P_{\theta}(e|e^{h}_{i},r_{i}),$
$\text{Acc}(\theta,G)=\sum_{i=1}^{M}$ $\mathbbm{1}[\hat{e}_{i}=e^{t}_{i}]/M,\qquad\ell(\theta,G)=\sum_{i=1}^{M}-\log P_{\theta}(e^{t}_{i}|e^{h}_{i},r_{i})/M.$
3 Real-world Experiments
Figure 1: The multiple-choice accuracy/loss on unseen triples of different-sized language models trained on a real-word knowledge graph FB15K-237. The left panel (trained with 10k steps) shows that the testing accuracy decreases after a certain model size. The middle panel shows U-shape loss curves of language models trained with different number steps. The right panel shows Note that the model size on x-axis is in log scale.
In our initial sets of experiments, we investigate the reasoning scaling effect using a real-world knowledge graph, FB15K-237 (Toutanova & Chen, 2015). FB15K-237 is sampled from FB15K (Bordes et al., 2013), which is a dataset adapted from the Freebase knowledge base (Bollacker et al., 2007), a web-scale knowledge base released by Google. FB15K-237 contains $N_{e}=14,505$ entities, $N_{r}=237$ relations, and $N=310,116$ knowledge triples.
In Figure 1, we show different-sized language models trained on FB15K-237 with different numbers of steps. When trained with the same number of steps, we observe a slight reasoning performance drop when using larger models. So we then look at the testing loss on these datasets and observe a U-shape trend with respect to the model size. This observation contradicts the previous belief that larger models yield a smaller testing loss. The training loss decreases monotonically with respect to model size.
This implies that a language model can overfit to the training data when it is overparameterized for the underlying reasoning structure. Such deviation from traditional scaling law has also been reported in broken neural scaling law (Caballero et al., 2023) which proposed a double-descent-like (Nakkiran et al., 2020) function form instead of a monotonic power-law form. There have also been observations of tasks with inverse scaling (Wei et al., 2023) for large language models.
In this paper, we mainly focus on the scaling of model size. Instead of only scaling the size of the training data, we explore different possible ways of generating the knowledge graph and studying the effect of overall graph complexity on the model reasoning performance. In the following sections, we will mostly focus on understanding the ”turning point” of the reasoning loss. More specifically, we want to understand what is the optimal model size that can obtain the smallest possible reasoning testing loss.
As shown in Figure 1 and in Section 5, we find this optimal model size is largely stable when training the model for enough steps. Note that, at training time, we repeat the training triples for many epochs (e.g. 30 times for FB15K-237) to find the optimal model size. This graph epoch is different from the real-world cases where we repeat the whole pretraining corpus for certain epochs. Because we can view each triple in the graph as a piece of factual knowledge (e.g. Barack Obama’s wife is Michelle Obama), this knowledge is usually repeated many times in a pretraining text corpus, in many different forms. Therefore, although our models have seen the same triple many times during training, the same piece of factual knowledge could also have been repeated several times in one pass of a real-world pretraining corpus.
4 Synthetic Data Construction
Figure 2: Nine possible node types generated by two logical rules. Each entity position in a rule would create a new entity type. Each relation shared between two rules would also create two new entity types.
To investigate how the underlying knowledge structure influences language models’ reasoning performance, we propose an algorithm to generate synthetic knowledge graphs that mimic real-world knowledge graphs. More specifically, we assume that the knowledge graph generation process is governed by a set of logical rules.
For example, a rule for inferring the locatedIn relation can be ($e_{1}$, locatedIn, $e_{2}$) $\leftarrow$ ($e_{1}$, neighborOf, $e_{3}$) $\wedge$ ($e_{3}$, locatedIn, $e_{2}$). Formally, for a target relation $r$, we consider logic rules with conjunctive form. For $\forall\{e_{i}\}_{i=0}^{n}\subset\mathcal{E}$,
$(e_{0},r,e_{n})\leftarrow(e_{0},r_{1},e_{1})\wedge...\wedge(e_{n-1},r_{n},e_{n}),$
where $(e_{i-1},r_{i},e_{i})\in\mathcal{G}$. We abbreviate such rule by $h(r)=[r_{1},r_{2},...,r_{n}]$. We randomly generate a set of logical rules $\mathcal{H}$ and ensuring there is no cycles in the set. To grow a graph that follows these rules, we enforce sparsity of the possible relation types connecting to and branching out each entity. More specifically, we define node types based on the possible relation types connecting to and branching out each entity, based on the generated rules, as illustrated in Figure 2. Such sparsity is also observed in real-world knowledge graphs.
Our random graph generation process is inspired by the preferential attachment process (Barabási & Albert, 1999), which is used for generating scale-free networks with a power-law distribution for the degrees of the nodes. Intuitively, preferential attachment implies a “the rich get richer” approach to edge placement in the graph. Each time a new node is added to the graph, there is a ‘preference’ to connect to the nodes that are already highly connected, with a probability proportional to the target node’s degree. Since we have observed the scale-free property in real-world knowledge graphs and the internet is known to be a scale-free network, we adopt a preferential attachment based graph generation process. To accommodate different relation types assigned to each edge, we maintain a degree distribution for each relationship and add new edges according to preferential attachment.
The code for our random graph generation algorithm is shown in the Appendix B. In summary, we first randomly generate a set of rules $\mathcal{H}$, with the number of rules $|\mathcal{H}|=N_{h}$ and the range of rule length $[L_{min},L_{max}]$ as hyperparameters. Then we generate all possible node types as illustrated in Figure 2, with the maximum number of relations per node $M_{r}$ as a hyperparameter. We generate a seed graph by instantiating each rule with a set of new entities. To this, we incrementally add one new entity until the number of entities reaches $N_{r}$, by first randomly assigning a node type to it, and then randomly sampling the $m$ relation types from the set of relations defined by the node type. We choose the target of these $m$ new edges by preferential attachment. After adding every $K$ entities, we search through the current graph to add any edges that can be inferred through the logic rules defined in $\mathcal{H}$. We call the triples that can be deduced through a logic rule by deductible triples, otherwise atomic triples.
Finally, we limit the number of training triples to $N$ and ensure that the the ratio between the number of deductible triples and atomic triples to $\gamma$ by subsampling the generated graph. We also further ensure that the triples in the held-out test set are all deductible through the training triple. In this way, we can generate synthetic knowledge graphs with specific sizes and complexity.
5 Scaling Laws
In this section, we investigate the scaling law of language models trained on different synthetic knowledge graphs. We conduct controlled experiments to show the effect of individual components of the data generation process. We also propose an information-theoretical way to measure the overall reasoning complexity of a knowledge graph, which we call the graph search entropy, and relate this linearly with the optimal model size. i.e. the model size that obtains the lowest possible testing loss.
5.1 Graph Generation Ablation
Figure 3: We show the effect of different hyperparameters of the synthetic knowledge graph generation process. In each experiment, we keep all other parameters the same and only change one hyperparameter. We show the effect with both the testing accuracy (left) and the testing loss (right) as the y-axis, with different model sizes as the x-axis in log scale.
We study the effects of the following four hyperparameters of graph data generation: the number of triples $N$, the number of entities $N_{e}$, the number of relations $N_{r}$, and the number of rules $N_{h}$. We fix all training hyperparameters as specified in the Appendix A but study the effect of training steps, as according to our preliminary experiments it has the largest effect on the optimal model size. The detailed data generation configuration for each set of experiments can also be found in the Appendix A.
Stable optimal model size with respect to training steps. In Figure 3 (a), we show the effect of training language models on the same knowledge graph with different numbers of training steps. As mention in the last part of Section 2, the optimal model size becomes smaller when the number of training steps increases, and then becomes stable after 4k steps. Another observation is regardless of the number of training steps, the maximum accuracy or minimum loss is stable. While we have ensured that all testing triples can be deduced through the training triples, there seems to be a performance cap determined solely by the knowledge graph data, which is unaffected by model size. In all following experiments, we train all models for 10k steps.
More triples implies a larger optimal model size. In Figure 3 (b), we show the effect of the number of unique triples $N$ sampled after the same knowledge graph generation process. This setting is arguably the most similar to the real-world pretraining of language models: the underlying world knowledge graph of all the pretraining corpora is largely stable, and training data are realizations of the underlying knowledge graph and so the sizes of different corpora are simply a result of subsampling/upsampling the knowledge in the existing graph. We can see that a larger number of training triples results in a larger optimal model size and a better reasoning performance. This observation aligns with the classic scaling laws. However, there exists an optimal model size for the full knowledge graph: after sampling beyond the size of the full knowledge graph, you can only sample previously seen knowledge. In this case, the optimal model size would be stable no matter the training data size.
Number of rules does not impact optimal model size. In Figure 3 (c), we show the effect of generating knowledge graphs of the same size with different numbers of rules $N_{h}$. More rules mean that the testing triples need to be solved in more ways. The number of rules does not have a significant effect on the optimal model size, but affects the reasoning performance. There appears to be an optimal number of rules (20) that results in the best performance. This is because more rules increases the complexity of solving the test set while fewer rules increases the ambiguity in the training set. i.e. a relation may be be deduced through correlations outside of the predefined rules. The reason why the number of rules does not affect the optimal model size is likely because it does not significantly impact the graph search entropy. This will be discussed in detail in Section 5.2.
More relations imply a larger optimal model size. In Figure 3 (d), we show the effect of generating knowledge graphs of the same size and the same number of rules with different numbers of relations $N_{r}$. While the rules used for deducing the testing set remain the same for all experiments, there are additional relations that may not be used by any of the rules. We construct knowledge graphs with an excessive number of relations by adding additional relation patterns. In general, more relations improves the best reasoning performance while increasing the optimal model size. More relations increases the complexity of the knowledge graph, and thus increases the optimal model size. On the other hand, as discussed in the previous experiment, a small number of rules along with a small number of relations increases the ambiguity in the training set. By adding dummy relations that are not used for reasoning, the language model can better distinguish between the logic rules and spurious correlations between relations. Thus the reasoning performance improves with more relations.
The optimal model size increases with the deductible ratio when the ratio is small. In Figure 3 (e), we show the effect of generating knowledge graphs with different ratios between deductible triples and atomic triples, $\gamma$, while keeping the number of entities and the number of triples unchanged. A larger ratio implies that the language model can see more rule patterns at training time, thus improving the reasoning performance. The increase in performance and optimal model size stops after a ratio threshold.
More entities imply a larger optimal model size. In Figure 3 (f), we show the effect of generating knowledge graphs with different numbers of nodes/entities $N_{e}$. In this experiment, we also scale the number of triples to keep all other hyperparameters unchanged. Increasing the number of entities increases the optimal model size while also increasing the testing loss. More entities imply a larger graph which increases the graph complexity, thus increasing the optimal model size. As in this experiment, we use a small number of rule ($N_{h}=5$) and relations ($N_{r}=10$), an excessive number of entities and triples will create more ambiguity thus hurting the reasoning performance.
5.2 Optimal Model Size v.s. Graph Search Entropy
Figure 4: The optimal model size with the lowest possible testing loss v.s. the graph search entropy. The red line is the linear regression line using data from the synthetic experiments (blue squares), with a 95% confidence interval. We also plot the graph search entropy and optimal model size from the real-world FB15K-237 experiment (green dot) to verify the accuracy of the obtained linear scaling law.
From our previous ablation studies, we hypothesize that the optimal model size is positively related to the overall complexity of the knowledge graph. Thus, we propose that we measure the complexity of a knowledge graph by quantifying the amount of information that can be obtained from the graph by exploring the graph through a random search. From our task definition, to reason over the knowledge graph, the language model needs to (a) identify the set of logic rules by observing repetitive patterns; (b) traverse the graph using one or more specific logic rules to locate the tail entity. So we define the graph search entropy as the maximum amount of information that can be obtained when randomly traversing the graph.
To simplify the problem, we first focus on the average amount of information we can observe at one node of the graph. If we consider a random walk over the knowledge graph, then we refer to the entropy produced by each step/node on the walk trace for an infinitely long random walk as the entropy rate of this random walk. For a graph $G$, the maximum entropy rate is equal to the log of the largest eigenvalue of the adjacency matrix $A$. Note that only consider the entropy rate with respect to the entity, without considering the entropy rate with respect to the relation. We can compute the relation entropy rate with the stationary distribution and transition matrix induced by the maximal entropy rate random walk. If we denote the dominating eigenvalue by $\lambda\in\mathbb{R}$ and the corresponding eigenvector by $\psi\in\mathbb{R}^{N_{e}}$, then the stationary distribution $\rho\in\mathbb{R}^{N_{e}}$ can be written as:
$\rho_{i}=\psi_{i}/||\psi||_{2}^{2}.$
The transition matrix $S\in\mathbb{R}^{N_{e}\times N_{e}}$ of the maximal entropy random walk can be written as:
$S_{ij}=(A_{ij}/\lambda)(\psi_{j}/\psi_{i}).$
We can then transform the entity-to-entity transition matrix $S\in\mathbb{R}^{N_{e}\times N_{e}}$ into an entity-to-relation transition matrix $S^{r}\in\mathbb{R}^{N_{e}\times N_{r}}$ by merging the entries with the same relation together:
$S^{r}_{ij}=\sum_{k=1}^{N_{e}}\mathbb{1}[(i,j,k)\in G]S_{ik}.$
Finally, the relation entropy rate $H^{r}(G)$ can be written as:
$H^{r}(G)=-\sum_{i=1}^{N_{e}}\rho_{i}\sum_{j=1}^{N_{r}}S^{r}_{ij}\log(S^{r}_{ij}).$
The overall graph search entropy $H(G)$ can then be written as the sum of the entity entropy rate and the relation entropy rate multiplied by the number of nodes:
$H(G)=N_{e}(\log(\lambda)+H^{r}(G)).$
We empirically investigate the relation between the optimal model and the graph search entropy by plotting them against each other in Figure 4, and perform linear regression. The optimal model sizes are obtained from the synthetic experiments conducted in the ablation studies. In the ablation studies we only report the results for exponentially increasing model sizes for clarity. In this study to better capture the optimal model size, we make the model sizes near the optimal model size more fine-grain. In all experiments, we keep the training hyperparameter the same, with 10k train steps.
We find a strong linear relation between the optimal model size and the graph search entropy with $R^{2}=0.85$. Note that there are a few sources of noise for locating the optimal model size for a specific knowledge graph. First, we only train language model with selected sizes due to compute and time limitations, and the quantization of the model size would disrupt the smoothness of the scaling law. Second, the exact location of the optimal model size is dependent on the training steps, which we did not thoroughly traverse but choose to inspect at the training step 10k.
After fitting a linear regression line using the data from our synthetic experiments, we check the validity of this empirical scaling law against our real-world knowledge graph, FB15K-237. We calculate the graph search entropy for FB15K-237, and find the predicted optimal model size is very close to the observed optimal model size, shown as a green dot in Figure 4.
From our scaling law, we can see that roughly 124 additional parameters in the optimal model size are required per 1-bit entropy increase in the knowledge graph. That is a language model can only reliably (not perfectly) reason over 0.008 bit information per parameter. This is very different from the knowledge capacity scaling law concluded by Allen-Zhu & Li (2025), which shows that the language model can store 2 bits of knowledge per parameter. We think this discrepancy is due to two reasons: first, our scaling law is not only about memorizing the knowledge, but also about reasoning over the learned knowledge, which is significantly harder. Second, the way we compute the graph search entropy is fundamentally different from the way Allen-Zhu & Li (2025) computes the knowledge entropy. While Allen-Zhu & Li (2025) describes the entropy of the knowledge generation process, our graph search entropy describes the entropy of randomly traversing a fixed knowledge graph. In this way, we did not directly measure the amount of information that a language model needs to memorize, but measuring the complexity of traversing, and therefore, reasoning over a graph. It is hard, if not impossible, to obtain the data generation process of real-world data, but it is possible to get an estimate of the underlying knowledge graph of a corpus through automated knowledge graph construction algorithms (Zhong et al., 2023). Thus, it is possible to predict the optimal reasoning model size for real-world pretraining, by first constructing a knowledge graph from the pretraining corpus, and then computing its graph search entropy, and finally using a similar scaling law to calculate the optimal model size.
5.3 Limitations
We want to highlight that this study is only conducted on simplified pretraining data from knowledge graphs, and the results are likely not directly applicable to real-world language model pretraining with large text corpus. The setting of our study provides a reasonable analogy to the real-world language model pretraining, and the obtained insight might be found useful in the real world when the compute is abundant with very large models and very large datasets that exhaustively traverse the underlying knowledge graph. We leave the work of verifying our scaling law in the real word to future research due to its resource-demanding nature.
6 Related Work
Language Model Scaling Laws
Kaplan et al. (2020) first observed a power-law relationship between LLM perplexity, model parameter count, and training data size, laying the foundation for scaling law research. Subsequently,  Hoffmann et al. (2022b) explored optimal training strategies under constrained computational resources and discovered that LLM parameter size and the number of training tokens should scale proportionally to achieve optimal compute efficiency under a fixed budget. Beyond pretraining performance, researchers further confirmed that downstream task performance can also be reliably predicted based on model size and training data volume (Hernandez et al., 2021; Isik et al., 2024).  Allen-Zhu & Li (2025); Lu et al. (2024) have turned to exploring more specific capability dimensions, focusing particularly on the scaling laws of factual memory in LLMs and their behavioral patterns when memorizing different types of facts. Most recently, Roberts et al. (2025) have confirmed that scaling laws are skill-dependent, and found that knowledge-intensive tasks are more parameter-hungry while reasoning-intensive tasks are more data-hungry.  Springer et al. (2025) challenge a core assumption in scaling research—that more pretraining invariably leads to better downstream performance. Our paper identifies a different U-shaped scaling curve under the specific scenario of knowledge graph reasoning and reveals that the search complexity of the knowledge graph determines the optimal model size. This echoes the discovery of Pandey (2024) and Yin et al. (2024) that classic scaling laws are highly dependent on the data complexity or the compression ratio of the data. Havrilla & Liao (2024) also confirmed from both theoretical and empirical perspectives that the power of the power scaling law depends on the intrinsic dimension of the training data.
Language Model Reasoning
Our paper focuses on the reasoning capability of language models which has drawn a lot of attention recently (Zhang et al., 2023; Chen et al., 2023; Yao et al., 2023a; b; Wang et al., 2023; Guo et al., 2025; Jin et al., 2024; Yeo et al., 2025; Team et al., 2025; Li et al., 2025). LLMs usually reason in a step-by-step manner in real-world tasks like math word problems (Wei et al., 2022b). In our experiments, we do not ask language models to generate a step-by-step solution for its answer, but ask the language model to directly choose the correct answer from the given options, because our pretrain-only language models are not trained to give a step-by-step solution for a query. Our synthetic reasoning environment is the most similar to Wang et al. (2024b), which also use the knowledge graph completion task as a testbed to understand how language models learn to reason at pretraining time. They propose that language models are able to aggregate random walk paths sampled from the knowledge graph. Wang et al. (2024a); Zhu et al. (2024) also employ a graph structure to ground their synthetic reasoning tasks to explain how LLMs reason, but their reasoning is defined as concatenations of relations: A is $r_{1}$ to B and B is $r_{2}$ to C implies A is $r_{1}r_{2}$ to C. The knowledge graph completion task we employ is more complex than simple concatenation of relations as the language model needs to find out which relation $r_{1}r_{2}$ corresponds to from the knowledge graph.
7 Conclusion
This paper investigates reasoning scaling in language models trained on knowledge graphs. Our results reveal a U-shaped relationship between model size and reasoning performance, where overparameterization leads to excessive memorization and degraded reasoning ability. We identify key factors that determine the optimal model size, such as the number of training triples and graph complexity. Notably, we propose an empirical scaling law linking optimal model size to graph search entropy, offering a quantitative guide for model design. While our experiments are conducted in controlled settings, these insights pave the way for future work in real-world pretraining scenarios and improved reasoning capabilities in LLMs.
Appendix A Experiment Details
Model size hidden size MLP size #attention heads #layers
0.3M 128 256 2 2
0.7M 128 256 2 4
1.3M 256 512 4 2
2.6M 256 512 4 4
5.3M 256 512 4 8
10.5M 512 1024 8 4
21.0M 512 1024 8 8
42.0M 512 1024 8 16
83.9M 1024 2048 16 8
167.8M 1024 2048 16 16
335.6M 1024 2048 16 32
671.2M 2048 4096 32 16
1342.4M 2048 4096 32 32
Table 1: Language model (Llama) size details
batch size lr lr scheduler warmup ratio weight decay max length
1024 1e-4 cosine 0.2 0 128
Table 2: Hyperparameter settings for language model pretraining.
$N$ $N_{e}$ $N_{r}$ $N_{h}$ $\gamma$
(a) 100k 10k 100 50 0.5
(b) 10k/20k/…/100k 10k 100 50 0.5
(c) 100k 10k 100 5/10/…/50 0.5
(d) 100k 10k 10/20/…/100 50 0.5
(e) 100k 10k 100 50 0.1/0.5/…/0.9
(f) 10k/20k/…/100k 1k/2k/…/10k 10 5 0.5
Table 3: Knowledge graph hyperparameter settings for Figure 3 experiments. We keep $L_{min}=2$ and $L_{max}=4$ for all experiments. Here $N$ denotes the number of triples, $N_{e}$ denotes the number of entities, $N_{r}$ denotes the number of relations, $N_{h}$ denotes the number of rules, $\gamma$ denotes the ratio between deductible triples and atomic triples, $L_{min}$ denotes the minimum rule length, and $L_{max}$ denotes the maximum rule length.
Appendix B Synthetic Knowledge Graph Generation Code
import networkx as nx
import numpy as np
import random
from collections import defaultdict
def add_edge(G, h, t, r):
num_edges = 0
if G.has_edge(h, t):
if r not in G[h][t][’id’]:
G[h][t][’id’].append(r)
num_edges += 1
else:
print(’edge already exists’)
else:
G.add_edge(h, t, id=[r])
num_edges += 1
print(’add edge: , (h, r, t), ’num edges: , num_edges)
return num_edges
def generate_rules(relations, num_rules, L_min, L_max, weighted=False, temperature=0.25):
# Generate K acyclic logic rules with varying lengths
dependency_graph = defaultdict(set)
rules = []
weights = []
if weighted:
for l in range(L_min, L_max + 1):
weights.append(np.exp(-temperature*l))
probs = np.array([w / sum(weights) for w in weights])
else:
weights = [1] * (L_max - L_min + 1)
def has_cycle(start, visited, stack):
”””Detects if adding a new dependency introduces a cycle.”””
if start not in visited:
visited.add(start)
stack.add(start)
print(’visited: , visited)
print(’stack: , stack)
for neighbor in dependency_graph[start]:
if neighbor in stack:
return True
elif has_cycle(neighbor, visited, stack):
return True
if start in stack:
stack.remove(start)
return False
for _ in range(num_rules):
while True:
if weighted:
length = random.choices(range(L_min, L_max + 1), weights=weights)[0]
else:
length = random.randint(L_min, L_max)
rule_relations = random.choices(relations, k = length + 1) # the first element is the implied relation
valid_rule = True
for i in range(1, len(rule_relations)):
dependency_graph[rule_relations[0]].add(rule_relations[i])
# Check for cycles
if has_cycle(rule_relations[i], set(), set()):
valid_rule = False
for j in range(1, i + 1):
dependency_graph[rule_relations[0]].remove(rule_relations[j])
break
if valid_rule:
rules.append(tuple(rule_relations))
break
print(’rules: , rules)
return rules
def get_node_types(rules, max_num_relations_per_node=3):
# map node types to out relations
node_types = {}
# map out relations to node types
r2node_types = defaultdict(list)
for rule in rules:
for i in range(len(rule)):
node_type = len(node_types)
if i == 0:
node_types[node_type] = [rule[i], rule[1]]
r2node_types[rule[i]].append(node_type)
r2node_types[rule[1]].append(node_type)
elif i == len(rule) - 1:
node_types[node_type] = [’-’ + rule[i], ’-’ + rule[0]]
r2node_types[’-’ + rule[i]].append(node_type)
r2node_types[’-’ + rule[0]].append(node_type)
else:
node_types[node_type] = [’-’ + rule[i], rule[i+1]]
r2node_types[’-’ + rule[i]].append(node_type)
r2node_types[rule[i+1]].append(node_type)
print(node_types)
print(r2node_types)
for num_rs in range(2, max_num_relations_per_node):
possible_new_node_types = []
for r in r2node_types:
alt_rs = []
for node_type in r2node_types[r]:
for _r in node_types[node_type]:
if _r != r:
alt_rs.append(_r)
alt_rs = list(set(alt_rs))
for node_type in r2node_types[r]:
if len(node_types[node_type]) == num_rs:
for _r in alt_rs:
if _r not in node_types[node_type]:
possible_new_node_types.append(tuple(sorted([_r] + list(node_types[node_type]))))
print(possible_new_node_types)
possible_new_node_types += list(set(possible_new_node_types))
possible_new_node_types = list(set(possible_new_node_types))
print(possible_new_node_types)
for rs in possible_new_node_types:
new_node_type = len(node_types)
node_types[new_node_type] = list(rs)
for _r in rs:
r2node_types[_r].append(new_node_type)
return node_types
def get_adj_out_relations(rules):
adj = defaultdict(list)
for rule in rules:
for i in range(len(rule)):
if i == 0:
adj[rule[i]].append(rule[1])
adj[rule[1]].append(rule[i])
elif i == len(rule) - 1:
adj[’-’ + rule[i]].append(’-’ + rule[0])
adj[’-’ + rule[0]].append(’-’ + rule[i])
else:
adj[’-’ + rule[i]].append(rule[i+1])
adj[rule[i+1]].append(’-’ + rule[i])
return adj
def latent_rule_graph(num_rules=50, L_min=2, L_max=4, n=10000, m=10, n_r=200,
num_test=1000, num_train=150000, check_frequency=100,
power_law=False, initial_graph=None,
length_weighted=False, mcmc=0.2, temperature=0.25,
deductible_ratio=0.5):
# Generate relations and entities
print(”mcmc: , mcmc)
relations = [’P’ + str(i) for i in range(n_r)]
all_rules = generate_rules(relations, max(n_r//L_min, num_rules), L_min, L_max)
r2rules = {}
for rule in all_rules:
if rule[0] not in r2rules:
r2rules[rule[0]] = []
r2rules[rule[0]].append(rule[1:])
num_triples = 0
repeated_entities = defaultdict(list) # map in relation to entities
child_relations = []
for rule in all_rules:
child_relations += rule[1:]
child_relations = list(set(child_relations))
child_relations += [’-’ + r for r in child_relations]
deductible_rules = random.sample(all_rules, num_rules)
if length_weighted:
weights = [int(100*np.exp(-temperature*len(rule))) for rule in all_rules]
else:
weights = [1 for _ in all_rules]
repeated_rules = []
for rule, weight in zip(all_rules, weights):
for _ in range(weight):
repeated_rules.append(rule)
random.shuffle(repeated_rules)
adj = get_adj_out_relations(repeated_rules)
all_deductibles = {}
if initial_graph is None:
# Default initial graph
G = nx.DiGraph()
node_id = 0
min_repeated_entities = 0
while min_repeated_entities < m:
for rule in all_rules:
source = ’Q’ + str(node_id)
node_id += 1
h = source
for r in rule[1:]:
t = ’Q’ + str(node_id)
node_id += 1
num_triples += add_edge(G, h, t, r)
repeated_entities[r].append(t)
repeated_entities[’-’ + r].append(h)
h = t
num_triples += add_edge(G, source, t, rule[0])
repeated_entities[rule[0]].append(t)
repeated_entities[’-’ + rule[0]].append(source)
min_repeated_entities = min([len(set(repeated_entities[r])) for r in child_relations])
else:
if len(initial_graph) < m or len(initial_graph) > n:
raise nx.NetworkXError(
f”Initial graph needs between m={m} and n={n} nodes”
G = initial_graph.copy()
node_id = len(G)
if not power_law:
repeated_entities = {r: list(set(repeated_entities[r])) for r in repeated_entities}
# Start adding the other nodes.
while node_id < n:
source = ’Q’ + str(node_id)
node_id += 1
possible_relations = [_r for _r in adj if _r in child_relations]
if len(possible_relations) == 0:
print(’no adj relations’)
break
print(’add child edge’)
chosen_edges = []
stop = False
for _ in range(m):
it = 0
while (r, t) in chosen_edges:
r = random.choice(possible_relations)
t = random.choice(repeated_entities[r])
it += 1
if it > 100:
print(’failed to find edge’)
stop = True
break
if stop or len(possible_relations) == 0:
break
possible_relations = [_r for _r in adj[r] if _r in child_relations]
chosen_edges.append((r, t))
if r[0] == ’-’:
num_triples += add_edge(G, t, source, r[1:])
repeated_entities[r[1:]].append(source)
else:
num_triples += add_edge(G, source, t, r)
repeated_entities[’-’ + r].append(source)
repeated_entities[r].append(t)
if len(possible_relations) == 0:
print(’no adj relations’)
break
if not power_law:
repeated_entities = {r: list(set(repeated_entities[r])) for r in repeated_entities}
if node_id % check_frequency == 0 or node_id == n-1:
# add deductibles
all_nodes = list(G.nodes)
random.shuffle(all_nodes)
for h in all_nodes:
for rule in deductible_rules:
head_list = [h]
r = rule[0]
for _r in rule[1:]:
next_head_list = []
for e_h in head_list:
if e_h not in G.nodes:
continue
for e_t in G[e_h]:
if _r in G[e_h][e_t][’id’]:
if random.random() < mcmc:
next_head_list.append(e_t)
head_list = next_head_list
for t in head_list:
if (h, r, t) not in all_deductibles:
all_deductibles[(h, r, t)] = [rule]
elif rule not in all_deductibles[(h, r, t)]:
all_deductibles[(h, r, t)].append(rule)
if not G.has_edge(h, t) or r not in G[h][t][’id’]:
print(’add deductible edge’)
add_edge(G, h, t, r)
num_triples += 1
repeated_entities[r].append(t)
repeated_entities[’-’ + r].append(h)
atomic_triples = []
deductible_triples = []
for h, t in G.edges:
for r in G[h][t][’id’]:
if (h, r, t) not in all_deductibles:
atomic_triples.append((h, r, t))
else:
deductible_triples.append((h, r, t))
random.shuffle(atomic_triples)
random.shuffle(deductible_triples)
assert len(atomic_triples) >= int(num_train * (1-deductible_ratio))
assert len(deductible_triples) >= int(num_train * deductible_ratio) + 2 * num_test
remove_triples = []
train_atomic_triples = atomic_triples[:int(num_train * (1-deductible_ratio))]
remove_triples += atomic_triples[int(num_train * (1-deductible_ratio)):]
train_deductible_triples = deductible_triples[:int(num_train * deductible_ratio)]
remove_triples += deductible_triples[int(num_train * deductible_ratio):]
for h, r, t in remove_triples:
_t = t
rs = G[h][_t][’id’]
if r in rs:
if len(rs) == 1:
G.remove_edge(h, _t)
else:
G[h][_t][’id’].remove(r)
train_triples = train_deductible_triples + train_atomic_triples
random.shuffle(train_triples)
print(”num train triples: , len(train_triples))
r2rule = {}
for rule in deductible_rules:
if rule[0] in r2rule:
r2rule[rule[0]].append(rule[1:])
else:
r2rule[rule[0]] = [rule[1:]]
def check_deductible(triple):
h, r, t = triple
alt_ts = []
for rule in r2rule[r]:
head_list = [h]
for _r in rule:
next_head_list = []
for e_h in head_list:
for e_t in G[e_h]:
if _r in G[e_h][e_t][’id’]:
next_head_list.append(e_t)
head_list = next_head_list
alt_ts += head_list
if t in alt_ts:
return True
return False
id_test_triples = []
for i in range(int(num_train * deductible_ratio), len(deductible_triples)):
if check_deductible(deductible_triples[i]):
id_test_triples.append(deductible_triples[i])
if len(id_test_triples) == num_test:
break
id_test_rules = [all_deductibles[triple] for triple in id_test_triples]
print(”num id test triples: , len(id_test_triples))
rule2triples = defaultdict(list)
for triple in deductible_triples[i+1:]:
for rule in all_deductibles[triple]:
rule2triples[rule].append(triple)
# uniformly sample testing triples from each rule
uniform_test_triples = []
for rule in rule2triples:
triples = []
for triple in rule2triples[rule]:
if check_deductible(triple):
triples.append(triple)
if len(triples) > num_test//len(rule2triples):
uniform_test_triples += random.sample(triples, num_test//len(rule2triples))
else:
uniform_test_triples += triples
random.shuffle(uniform_test_triples)
uniform_test_rules = [all_deductibles[triple] for triple in uniform_test_triples]
print(”num uniform test triples: , len(uniform_test_triples))
return G, deductible_rules, train_triples, id_test_triples, id_test_rules, uniform_test_triples, uniform_test_rules

### v3

@@
-Do Larger Language Models Imply Better Generalization? A Pretraining Scaling Law for Implicit Reasoning
+Do Larger Language Models Generalize Better? A Scaling Law for Implicit Reasoning at Pretraining Time
 Abstract
-Large Language Models (LLMs) have demonstrated remarkable capabilities across a wide range of tasks requiring complex reasoning.
-However, the effects of scaling on their reasoning abilities remain insufficiently understood.
-In this paper, we introduce a synthetic multihop reasoning environment designed to closely replicate the structure and distribution of real-world large-scale knowledge graphs.
-Our reasoning task involves completing missing edges in the graph, which requires advanced multi-hop reasoning and mimics real-world reasoning scenarios.
-To evaluate this, we pretrain language models (LMs) from scratch solely on triples from the incomplete graph and assess their ability to infer the missing edges.
-Interestingly, we observe that overparameterization can impair reasoning performance due to excessive memorization.
-We investigate different factors that affect this U-shaped loss curve, including graph structure, model size, and training steps.
-To predict the optimal model size for a specific knowledge graph, we find an empirical scaling that linearly maps the knowledge graph search entropy to the optimal model size.
-This work provides new insights into the relationship between scaling and reasoning in LLMs, shedding light on possible ways to optimize their performance for reasoning tasks.
+Reasoning is an integral part of many tasks performed by language models (LMs).
+However, the effects of scaling model sizes and data on reasoning abilities at pretraining time remain understudied.
+To rigorously investigate this problem, we pretrain LMs from scratch on a synthetic implicit multihop reasoning environment designed to closely replicate the structure and distribution of real-world large-scale knowledge graphs.
+We then assess the LMs’ ability to complete the missing edges in the graph, which requires multi-hop reasoning that can be viewed as a simplification of implicit reasoning during real-world pretraining.
+Interestingly, we observe that overparameterization can impair the implicit reasoning performance due to excessive memorization.
+We investigate different factors that affect the loss curve when scaling different components of the knowledge graph, model size, and training steps.
+To predict the optimal model size for a specific knowledge graph, we find an empirical scaling law that shows optimal-sized LMs can approximately reason over 0.008 bit information per parameter.
+This work shows counterintuitive effects of model size scaling and provides new insights into the relationship between scaling and reasoning in LLMs.
+[^1] Code: https://github.com/WANGXinyiLinda/reasoning-scaling-law
 1 Introduction
-Large Language Models (LLMs) have demonstrated remarkable capabilities across a wide range of tasks.
-Recently, the reasoning capacity of LLMs has drawn a lot of attention as it is highly correlated with LLMs’ performance on many complex real-world tasks (Wei et al., 2022a; Guo et al., 2025).
-While the reasoning capability is usually enhanced during the post-training stage, it is reasonable to assume that LLMs have already acquired the capability during the pretraining stage, as the post-training is on a significantly smaller scale than pretraining.
-Some recent work has investigated the possible mechanism of LLMs learning to reason through next-token prediction pretraining (Zhu et al., 2024; Wang et al., 2024a; b).
-However, the impact of pretraining scaling on LLMs’ reasoning ability remains insufficiently understood.
-The general scaling behavior of language models has been investigated, including the well-known exponential scaling laws for testing loss and compute proposed by Kaplan et al.
+Language Models (LMs) have demonstrated remarkable capabilities across a wide range of tasks, with reasoning being a core component (Wei et al., 2022a; Guo et al., 2025).
+While reasoning is typically enhanced during the post-training stage by encouraging LMs to generate long chain-of-thoughts (CoTs) (Guo et al., 2025; Yang et al., 2025), it is reasonable to assume that they already acquire the foundations of such capability during pretraining, given that post-training operates at a significantly smaller scale.
+Several recent studies have explored the mechanisms by which LMs may acquire reasoning-related abilities through next-token prediction pretraining (Zhu et al., 2024; Wang et al., 2024a; b).
+However, the impact of scaling on LMs’ reasoning ability during pretraining remains poorly understood.
+The general scaling behavior of LMs at pretraining time has been extensively investigated, including the well-known exponential scaling laws for testing loss and compute proposed by Kaplan et al.
 (2020) and the training compute-optimal scaling studied by Hoffmann et al.
@@
 Recent work has also examined the scaling of specific capabilities like machine translation (Ghorbani et al., 2022) and knowledge capacity/memorization (Allen-Zhu & Li, 2025; Lu et al., 2024).
-According to these existing scaling laws, it is in general believed that larger models imply better testing loss or task performance when trained on more data.
-In this paper, however, we find that the reasoning ability of language models under pretraining scaling can behave differently from normal power-law scaling, in a simplified pretraining environment.
-More specifically, when given enough compute, we find that the testing loss scaling curve is U-shape and there exists an optimal model size that produces the best reasoning performance/testing loss.
-This implies that overparameterization might hurt reasoning capability during pre-training.
-We first observe this phenomenon with real-world knowledge graph data, and then systematically study it through synthetically generated data.
-We choose to mimic the real-world knowledge structure and distribution with synthetic knowledge graphs (KGs).
-We define reasoning over world knowledge as completing missing edges in an incomplete knowledge graph, which requires multiple jumps on the knowledge graph according to some pre-defined rules that are latently encoded into the graph generation process.
-To analyze this, we pretrain LMs from scratch using only triples from the incomplete graph and evaluate their ability to infer missing connections.
-We investigate important factors that affect the U-shape scaling of reasoning loss versus language model size. Our important findings can be summarized as follows:
-  • The minimum reasoning loss/maximum reasoning accuracy that a language model can reach is capped by the training data, regardless of the training steps and model size.
-  • The optimal model size for a training corpus is largely fixed regardless of the training steps when the number of training steps is large enough.
-  • When the underlying knowledge graph is fixed, training on more data sampled from the graph increases the optimal model size and reasoning performance.
-  • More complex knowledge graph implies a larger optimal model size.
-As we observed that the optimal model size is likely solely determined by the training knowledge graph, we then aim to find an empirical scaling law that can predict the optimal model size with knowledge graph statistics.
-We then discover a linear relationship between the optimal model size and a newly proposed graph search entropy, which measures the entropy of performing random searches on a knowledge graph.
-Roughly, 124 additional parameters in the optimal model size are required per 1-bit entropy increase in searching a knowledge graph.
-Our work contributes to the broader understanding of LLM reasoning by shedding light on the intricate relationship between scaling and reasoning capability.
-Our proposed empirical reasoning scaling law provides possible practical insights for optimizing LLMs’ reasoning ability at pretraining time.
-2 Preliminaries
-While the real-world LLMs are pretrained on large scale text corpus, this corpus can be viewed as encoding a wide range of world knowledge.
+According to these existing scaling laws, it is in general believed that larger models imply better testing loss or task performance.
+In this paper, instead we find that the scaling of LMs’ reasoning capability at pretraining time differs from normal power-law scaling, in a simplified controlled pretraining environment.
+We use implicit reasoning to denote the reasoning behavior that naturally emerges during pretraining.
+i.e.
+the capability to draw new conclusions from existing knowledge without being explicitly trained to generate chain-of-thoughts (CoTs).
+More specifically, we define implicit reasoning over world knowledge as the task of completing missing edges in an incomplete knowledge graph, which requires multi-hop traversal according to predefined logic rules that are implicitly encoded in the graph generation process.
+To investigate this, we pretrain LMs from scratch using only triples from the incomplete graph and then evaluate their ability to infer the missing connections.
+With sufficient compute, we find that the curve of implicit reasoning loss versus model size follows a U-shape, revealing an optimal model size that yields the best reasoning performance.
+This suggests that overparameterization may impair the implicit reasoning capability instilled during pretraining.
+We first observe this phenomenon using data derived from real-world knowledge graphs, and then systematically study it with synthetically generated data.
+We investigate important factors that affect the U-shaped scaling of reasoning loss versus language model size. Our important findings can be summarized as follows:
+  • The minimum reasoning loss reachable by an LM is solely determined by the training data, regardless of training steps and model size.
+  • The optimal model size is solely determined by knowledge graph complexity and data size regardless of training steps.
+  • We show that an optimal-sized LM can approximately reason over 0.008 bit information per parameter.
+As we observed that the optimal model size is likely solely determined by the training knowledge graph, we then aim to find an empirical scaling law that can predict the optimal model size from knowledge graph statistics.
+We identify a linear relationship between the optimal model size and our proposed graph search entropy, which quantifies the entropy of performing random searches on a knowledge graph.
+Under this framework, we find that each parameter in the optimal model size can reason over approximately 0.008 bits of information in a knowledge graph.
+In contrast, Allen-Zhu & Li (2025) show that a language model can memorize up to 2 bits of information per parameter—substantially more than its reasoning capacity.
+This gap arises both from the greater difficulty of reasoning compared to memorization and from the different methodologies used to compute these information quantities.
+A more detailed discussion is provided in Section 5.2.
+Our work contributes to the broader understanding of LLM reasoning by shedding light on the intricate relationship between scaling and implicit reasoning capability.
+Our proposed empirical reasoning scaling law provides possible practical insights for optimizing LLMs’ implicit reasoning ability at pretraining time.
+2 Method
+While real-world LLMs are pretrained on large scale text corpora, this corpus can be viewed as encoding a wide range of world knowledge.
 The power of LLMs lies in the fact that they can not only memorize the world knowledge and extract the knowledge when queried, but also reason over the world knowledge and draw novel conclusions.
-In this paper, we propose to construct a simplified pretraining corpus directly from a knowledge graph.
-A knowledge graph is comprised of a set of triples, and we use each knowledge triple as a training example.
+In this paper, we propose constructing a simplified pretraining corpus from a knowledge graph.
+A knowledge graph is comprised of a set of (head entity, relation, tail entity) triples, and we use each knowledge triple as a training example.
 We test the reasoning capability of a language model trained on such a corpus by testing its accuracy in completing triples that have never been seen in the knowledge graph but can be deduced through latent rules encoded in the graph structure.
@@
 We use the Llama (Touvron et al., 2023) model architecture to implement LMs of different sizes by adjusting the hidden dimensions and the number of layers.
-The specific parameter scheme can be found in the Appendix A.
-To evaluate the language model’s capability of reasoning over the knowledge graph, we test the language models on a held-out set of triples that are not seen in the training time.
+The specific parameter scheme can be found in the Appendix B.
+To evaluate the language model’s capability of reasoning over the knowledge graph, we test the LMs on a held-out set of triples that are not seen in the training time.
 Note that all entity and relation types should have been seen during training time and the language model is only tasked to connect missing edges.
@@
 $\text{Acc}(\theta,G)=\sum_{i=1}^{M}$ $\mathbbm{1}[\hat{e}_{i}=e^{t}_{i}]/M,\qquad\ell(\theta,G)=\sum_{i=1}^{M}-\log P_{\theta}(e^{t}_{i}|e^{h}_{i},r_{i})/M.$
-3 Real-world Experiments
-Figure 1: The multiple-choice accuracy/loss on unseen triples of different-sized language models trained on a real-word knowledge graph FB15K-237.
-The left panel (trained with 10k steps) shows that the testing accuracy decreases after a certain model size.
-The middle panel shows U-shape loss curves of language models trained with different number steps.
-The right panel shows Note that the model size on x-axis is in log scale.
+3 Initial Experiments with Real-World Knowledge Graph
+Figure 1: The multiple-choice accuracy/loss on unseen triples of different-sized LMs trained on a real-word knowledge graph FB15K-237.
+The first column shows that the testing accuracy decreases after a certain model size.
+The second column shows U-shape loss curves of LMs trained with different numbers of steps.
+The third column shows the training loss decreases steadily.
+These trends are stable across different ways of processing the knowledge triples, with the triple-only data shows the cleanest trend.
+Note that the model size on x-axis is in log scale.
 In our initial sets of experiments, we investigate the reasoning scaling effect using a real-world knowledge graph, FB15K-237 (Toutanova & Chen, 2015).
@@
 FB15K-237 contains $N_{e}=14,505$ entities, $N_{r}=237$ relations, and $N=310,116$ knowledge triples.
-In Figure 1, we show different-sized language models trained on FB15K-237 with different numbers of steps.
-When trained with the same number of steps, we observe a slight reasoning performance drop when using larger models.
-So we then look at the testing loss on these datasets and observe a U-shape trend with respect to the model size.
-This observation contradicts the previous belief that larger models yield a smaller testing loss.
-The training loss decreases monotonically with respect to model size.
+We process this dataset in three different ways: (a) translate each knowledge triple into a natural language sentence by prompting GPT4 and then tokenize the sentence with a pre-trained tokenizer, as shown in the first row of Figure 1; (b) translate each knowledge triple into a natural language sentence using pre-generated templates, as show in the second row of Figure 1; (c) translate each knowledge triple into text by assigning a random ID to each entity and relation and tokenize them by characters, as shown in the last row of Figure 1.
+An example can be found in Appendix A Figure 5.
+In Figure 1, we show different-sized LMs trained on FB15K-237 in all settings with different numbers of training steps.
+We observe a consistant reasoning performance drop when using larger models, across different ways of processing the knowledge triples, while the training loss decreases monotonically with respect to model size.
+This observation contradicts the previous belief that larger models always yield a smaller testing loss.
 This implies that a language model can overfit to the training data when it is overparameterized for the underlying reasoning structure.
 Such deviation from traditional scaling law has also been reported in broken neural scaling law (Caballero et al., 2023) which proposed a double-descent-like (Nakkiran et al., 2020) function form instead of a monotonic power-law form.
-There have also been observations of tasks with inverse scaling (Wei et al., 2023) for large language models.
-In this paper, we mainly focus on the scaling of model size.
-Instead of only scaling the size of the training data, we explore different possible ways of generating the knowledge graph and studying the effect of overall graph complexity on the model reasoning performance.
+There have also been observations of tasks with inverse scaling (Wei et al., 2023) for large LMs.
+In this paper, we focus primarily on the scaling of model size and data complexity.
+Rather than merely increasing the size of the training data, we explore many different setting for generating synthetic knowledge graphs.
+This allows us to ablate individual components of the graph generation process and examine how overall graph complexity affects reasoning.
+In the synthetic experiments presented below, we use random IDs instead of natural language sentences to eliminate lexical and syntactic effects, yielding cleaner trends from which we can draw quantitative conclusions.
 In the following sections, we will mostly focus on understanding the ”turning point” of the reasoning loss.
-More specifically, we want to understand what is the optimal model size that can obtain the smallest possible reasoning testing loss.
-As shown in Figure 1 and in Section 5, we find this optimal model size is largely stable when training the model for enough steps.
+More specifically, we want to understand what is the optimal model size, that is the model size that can obtain the smallest possible reasoning testing loss.
+As shown in Figure 1 and in Figure 3 (a), we find this optimal model size is largely stable when training the model for enough steps.
 Note that, at training time, we repeat the training triples for many epochs (e.g.
@@
 Each relation shared between two rules would also create two new entity types.
-To investigate how the underlying knowledge structure influences language models’ reasoning performance, we propose an algorithm to generate synthetic knowledge graphs that mimic real-world knowledge graphs.
+To investigate how the underlying knowledge structure influences LMs’ reasoning performance, we propose an algorithm to generate synthetic knowledge graphs that mimic real-world knowledge graphs.
 More specifically, we assume that the knowledge graph generation process is governed by a set of logical rules.
@@
 We abbreviate such rule by $h(r)=[r_{1},r_{2},...,r_{n}]$.
-We randomly generate a set of logical rules $\mathcal{H}$ and ensuring there is no cycles in the set.
+We randomly generate a set of logical rules $\mathcal{H}$ and ensure there are no cycles in the set.
 To grow a graph that follows these rules, we enforce sparsity of the possible relation types connecting to and branching out each entity.
@@
 Each time a new node is added to the graph, there is a ‘preference’ to connect to the nodes that are already highly connected, with a probability proportional to the target node’s degree.
-Since we have observed the scale-free property in real-world knowledge graphs and the internet is known to be a scale-free network, we adopt a preferential attachment based graph generation process.
+Since we have observed the scale-free property in real-world knowledge graphs and the internet is known to be a scale-free network (Albert et al., 1999), we adopt a preferential attachment based graph generation process.
 To accommodate different relation types assigned to each edge, we maintain a degree distribution for each relationship and add new edges according to preferential attachment.
-The code for our random graph generation algorithm is shown in the Appendix B.
-In summary, we first randomly generate a set of rules $\mathcal{H}$, with the number of rules $|\mathcal{H}|=N_{h}$ and the range of rule length $[L_{min},L_{max}]$ as hyperparameters.
+A comparison of the node degree distribution between synthetic graph and real-world graph can be found in Appendix C Figure 6.
+The code for our random graph generation algorithm is shown in the Appendix D.
+In summary, we first randomly generate a set of rules $\mathcal{H}$, with the number of rules $|\mathcal{H}|=N_{h}$ and the range of rule length $L_{min},L_{max}]$ as hyperparameters.
 Then we generate all possible node types as illustrated in Figure 2, with the maximum number of relations per node $M_{r}$ as a hyperparameter.
@@
 After adding every $K$ entities, we search through the current graph to add any edges that can be inferred through the logic rules defined in $\mathcal{H}$.
-We call the triples that can be deduced through a logic rule by deductible triples, otherwise atomic triples.
+We call the triples that can be deduced through a logic rule by deducible triples, otherwise atomic triples.
 Finally, we limit the number of training triples to $N$ and ensure that the the ratio between the number of deductible triples and atomic triples to $\gamma$ by subsampling the generated graph.
@@
 We study the effects of the following four hyperparameters of graph data generation: the number of triples $N$, the number of entities $N_{e}$, the number of relations $N_{r}$, and the number of rules $N_{h}$.
-We fix all training hyperparameters as specified in the Appendix A but study the effect of training steps, as according to our preliminary experiments it has the largest effect on the optimal model size.
-The detailed data generation configuration for each set of experiments can also be found in the Appendix A.
+We fix all training hyperparameters as specified in the Appendix B.
+In all experiments except Figure 3 (a), we train all models for 10k steps.
+The detailed data generation configuration for each set of experiments can also be found in the Appendix B.
 Stable optimal model size with respect to training steps.
 In Figure 3 (a), we show the effect of training language models on the same knowledge graph with different numbers of training steps.
-As mention in the last part of Section 2, the optimal model size becomes smaller when the number of training steps increases, and then becomes stable after 4k steps.
+As mentioned in Section 3, the optimal model size becomes smaller when the number of training steps increases, and then becomes stable after 4k steps.
 Another observation is regardless of the number of training steps, the maximum accuracy or minimum loss is stable.
 While we have ensured that all testing triples can be deduced through the training triples, there seems to be a performance cap determined solely by the knowledge graph data, which is unaffected by model size.
-In all following experiments, we train all models for 10k steps.
 More triples implies a larger optimal model size.
@@
 If we denote the dominating eigenvalue by $\lambda\in\mathbb{R}$ and the corresponding eigenvector by $\psi\in\mathbb{R}^{N_{e}}$, then the stationary distribution $\rho\in\mathbb{R}^{N_{e}}$ can be written as:
-$\rho_{i}=\psi_{i}/||\psi||_{2}^{2}.$
+[…]
 The transition matrix $S\in\mathbb{R}^{N_{e}\times N_{e}}$ of the maximal entropy random walk can be written as:
@@
 We can then transform the entity-to-entity transition matrix $S\in\mathbb{R}^{N_{e}\times N_{e}}$ into an entity-to-relation transition matrix $S^{r}\in\mathbb{R}^{N_{e}\times N_{r}}$ by merging the entries with the same relation together:
-$S^{r}_{ij}=\sum_{k=1}^{N_{e}}\mathbb{1}[(i,j,k)\in G]S_{ik}.$
+$S^{r}_{ij}=\sum_{k=1}^{N_{e}}\mathds{1}[(i,j,k)\in G]S_{ik}.$
 Finally, the relation entropy rate $H^{r}(G)$ can be written as:
@@
 5.3 Limitations
-We want to highlight that this study is only conducted on simplified pretraining data from knowledge graphs, and the results are likely not directly applicable to real-world language model pretraining with large text corpus.
+We want to highlight that this study is only conducted on simplified pretraining data from knowledge graphs, and the results are not directly applicable to real-world language model pretraining with large text corpus.
 The setting of our study provides a reasonable analogy to the real-world language model pretraining, and the obtained insight might be found useful in the real world when the compute is abundant with very large models and very large datasets that exhaustively traverse the underlying knowledge graph.
@@
 Language Model Reasoning
-Our paper focuses on the reasoning capability of language models which has drawn a lot of attention recently (Zhang et al., 2023; Chen et al., 2023; Yao et al., 2023a; b; Wang et al., 2023; Guo et al., 2025; Jin et al., 2024; Yeo et al., 2025; Team et al., 2025; Li et al., 2025).
-LLMs usually reason in a step-by-step manner in real-world tasks like math word problems (Wei et al., 2022b).
-In our experiments, we do not ask language models to generate a step-by-step solution for its answer, but ask the language model to directly choose the correct answer from the given options, because our pretrain-only language models are not trained to give a step-by-step solution for a query.
+Our paper focuses on the reasoning capability of LMs which has drawn a lot of attention recently (Zhang et al., 2023; Chen et al., 2023; Yao et al., 2023a; b; Wang et al., 2023; Guo et al., 2025; Jin et al., 2024; Yeo et al., 2025; Team et al., 2025; Li et al., 2025).
+LLMs are usually trained to reason in a step-by-step manner in real-world tasks like math problems (Wei et al., 2022b) and coding (Yang et al., 2024).
+In our experiments, we do not ask LMs to generate a CoT solution, but ask the language model to directly choose the correct answer from the given options, because our pretrain-only LMs are not trained to give a CoT solution for a query.
 Our synthetic reasoning environment is the most similar to Wang et al.
-(2024b), which also use the knowledge graph completion task as a testbed to understand how language models learn to reason at pretraining time.
-They propose that language models are able to aggregate random walk paths sampled from the knowledge graph.
+(2024b), which also use the knowledge graph completion task as a testbed to understand how LMs learn to reason at pretraining time.
+They propose that LMs are able to aggregate random walk paths sampled from the knowledge graph.
 Wang et al.
@@
 7 Conclusion
-This paper investigates reasoning scaling in language models trained on knowledge graphs.
-Our results reveal a U-shaped relationship between model size and reasoning performance, where overparameterization leads to excessive memorization and degraded reasoning ability.
-We identify key factors that determine the optimal model size, such as the number of training triples and graph complexity.
-Notably, we propose an empirical scaling law linking optimal model size to graph search entropy, offering a quantitative guide for model design.
-While our experiments are conducted in controlled settings, these insights pave the way for future work in real-world pretraining scenarios and improved reasoning capabilities in LLMs.
-Appendix A Experiment Details
+This paper presents a rigorous study of the scaling behavior of implicit reasoning in language models pretrained on knowledge graphs.
+Our findings reveal a U-shaped relationship between implicit reasoning performance and model size: overparameterization induces excessive memorization, which in turn degrades reasoning ability.
+We further identify key factors that determine the optimal model size, including the number of training triples and the complexity of the graph.
+Most notably, we propose an empirical scaling law that links the optimal model size to graph search entropy, demonstrating that a language model can reason over approximately 0.008 bits of information per parameter.
+Although our experiments are conducted in controlled settings to ensure rigor, the insights derived from this work offer promising directions for future studies on real-world pretraining and the enhancement of reasoning capabilities in large language models.
+8 Reproducibility statement
+We have taken several steps to ensure the reproducibility of our results.
+All experimental settings, including model architectures, training procedures, and hyperparameters, are described in detail in Section 2, Section 5 and Appendix B.
+To facilitate empirical reproducibility, we include a script of data construction steps in Appendix D.
+Additionally, we provide a simplified version of our source code as an easy-to-run Jupyter notebook for reproducing some of our experiments in the supplementary materials.
+Appendix
+Appendix A Data Processing Example
+Figure 5: An example of a triple being processed in three different ways.
+Appendix B Experiment Details
 Model size hidden size MLP size #attention heads #layers
@@
 Table 2: Hyperparameter settings for language model pretraining.
-$N$ $N_{e}$ $N_{r}$ $N_{h}$ $\gamma$
+NN $N_{e}$ $N_{r}$ $N_{h}$ $\gamma$
 (a) 100k 10k 100 50 0.5
@@
 We keep $L_{min}=2$ and $L_{max}=4$ for all experiments.
-Here $N$ denotes the number of triples, $N_{e}$ denotes the number of entities, $N_{r}$ denotes the number of relations, $N_{h}$ denotes the number of rules, $\gamma$ denotes the ratio between deductible triples and atomic triples, $L_{min}$ denotes the minimum rule length, and $L_{max}$ denotes the maximum rule length.
-Appendix B Synthetic Knowledge Graph Generation Code
+Here NN denotes the number of triples, $N_{e}$ denotes the number of entities, $N_{r}$ denotes the number of relations, $N_{h}$ denotes the number of rules, $\gamma$ denotes the ratio between deductible triples and atomic triples, $L_{min}$ denotes the minimum rule length, and $L_{max}$ denotes the maximum rule length.
+Appendix C Synthetic Knowledge Graph v.s. Real-world Knowledge Graph
+Figure 6: Distribution of node degrees of synthetic and real-world knowledge graphs.
+Figure 7: Distribution of number of outgoing relations per node of synthetic and real-world knowledge graphs.
+Appendix D Synthetic Knowledge Graph Generation Code
 import networkx as nx